# SentenceKV: Efficient LLM Inference via Sentence-Level Semantic KV Caching

### v1

@@ 5.4 Efficiency
Using the Llama-3.1-8B-Instruct model, SentenceKV achieves substantial memory savings as context length increases. At 256k tokens, it uses 52.71 GB of memory, compared to 89.71 GB for Full KV. Its inference latency remains stable at around 17.8 ms, significantly lower than Full KV’s 84.9 ms. Overall, SentenceKV supports longer contexts under constrained GPU memory while maintaining low latency, making it a practical and efficient solution for long-context decoding. See Appendix A.3 for results of the LongChat-v1.5-7B model.
@@ 6.1 Sentence chunking
An alternative strategy to our sentence-level chunking is to perform equal-sized chunking based on the total number of sentences—dividing the text uniformly. However, this approach consistently yields worse results on LongBench (see Table 2), likely due to the loss of coherence and topical alignment within chunks. When sentences are split arbitrarily, important semantic relationships and narrative flow can be disrupted, making it harder for the model to maintain context and generate meaningful outputs. In contrast, our approach leverages semantic segmentation to group related information together, preserving logical boundaries and thematic continuity. This experiment demonstrates that chunking based on semantic structure leads to better contextual integrity and improved downstream performance.
@@ 7 Conclusion
In this work, we presented SentenceKV, a novel sentence-level semantic KV caching framework that significantly enhances inference efficiency for large language models operating on long contexts. Our approach seamlessly integrates token-level and sentence-level operations by embedding sentence-level processing within the core token-level computations of transformer-based models. By grouping tokens into semantically coherent sentences, SentenceKV compresses the KV cache during prefilling into compact sentence embeddings on GPU while offloading detailed KV pairs to CPU. During decoding, only the most semantically relevant sentences are retrieved to satisfy a fixed token budget, preserving contextual coherence and minimizing redundant data transfers.
Extensive evaluations on the PG19, LongBench, and NeedleInAHaystack benchmarks demonstrate that SentenceKV matches full KV cache performance in perplexity and task accuracy - even using as little as 3% of the full cache capacity - while reducing GPU memory usage by more than 30% and maintaining stable low latency throughout context lengths up to 256K tokens. In retrieval-heavy settings, SentenceKV outperforms existing state-of-the-art methods by more than 20% in retrieval accuracy, highlighting its ability to locate and recall critical information within lengthy inputs.
@@
Despite its effectiveness, our approach exhibits certain limitations that future research should address. Firstly, our current reliance on punctuation marks for sentence segmentation may not always ensure accurate sentence boundaries, particularly in texts lacking clear punctuation. Such inaccuracies can degrade semantic coherence. Advanced natural language processing techniques, such as sentence segmentation tools provided by the Natural Language Toolkit [3], could significantly enhance accuracy in punctuation-deficient contexts. Secondly, the fixed semantic factor employed in our method may limit adaptability across texts with varying sentence lengths and complexities. Future research should explore adaptive methods that dynamically adjust the semantic factor based on sentence length distributions, aiming to optimize semantic coherence and processing efficiency across diverse textual inputs.
Appendix A Appendix / supplemental material
@@ A.2 SentenceKV: pseudocode
0:  Prompt tokens, token budget $\tau$, semantic keeping factor $r$, observation window size $N$
@@
9:  Decoding Phase:
10:  Initialize an empty sentence cache $Q_{s}$
11:  for $t=1,2,\dots$ do
12:     Generate next token $x_{t}$ and compute query $q_{t}$
13:     Append $q_{t}$ to $Q_{s}$
14:     if a sentence boundary is reached then
@@
20:        Reset $Q_{s}$ for the next sentence
21:     end if
22:  end for
@@ A.3 Latency and Memory Footprint
For the Longchat-v1.5-7B model, SentenceKV similarly demonstrates considerable efficiency gains. The peak memory footprint at the maximum context length (256k) is reduced from the OOM (out-of-memory) scenario encountered by Full KV caching to a manageable 48.31 GB. Latency also remains stable around 16.9 ms, whereas Full KV caching becomes infeasible beyond 64k tokens due to excessive memory demands. Results details are shown in Appendix Table 3 and Table 4.

### v2

@@ 5.5 Efficiency
Using the Llama-3.1-8B-Instruct model, SentenceKV achieves substantial memory savings as context length increases. At 256k tokens, it uses 52.71 GB of memory, compared to 89.71 GB for Full KV. Its inference latency remains stable at around 17.8 ms, significantly lower than Full KV’s 84.9 ms. Overall, SentenceKV supports longer contexts under constrained GPU memory while maintaining low latency, making it a practical and efficient solution for long-context decoding. See Appendix A.3 for results of the LongChat-v1.5-7B model. A detailed efficiency and accuracy tradeoff analysis is shown in Appendix A.5, and a detailed GPU profilling is shown in Appendix A.6.
@@ 6.1 Sentence chunking
An alternative strategy to our sentence-level chunking is to perform equal-sized chunking based on the total number of sentences, dividing the text uniformly. However, this approach consistently yields worse results on LongBench (see Table 3), likely due to the loss of coherence and topical alignment within chunks. When sentences are split arbitrarily, important semantic relationships and narrative flow can be disrupted, making it harder for the model to maintain context and generate meaningful outputs.
@@
In contrast, our approach leverages semantic segmentation to group related information together, preserving logical boundaries and thematic continuity. This experiment demonstrates that chunking based on semantic structure leads to better contextual integrity and improved downstream performance.
@@ 7 Conclusion
In this work, we introduced SentenceKV, a sentence‑level semantic KV caching framework that improves inference efficiency for large language models on long contexts. SentenceKV integrates sentence‑level processing into token‑level transformer computations by grouping tokens into coherent sentences. It compresses the KV cache into compact sentence embeddings on GPU and offloads full KV pairs to CPU. During decoding, only the most relevant sentences are retrieved within a fixed token budget, preserving coherence and reducing data transfer. Evaluations on PG19, LongBench, NIAH and RULER show that SentenceKV matches full-cache performance in perplexity and task accuracy using as little as 3% of the cache, while reducing GPU memory usage by over 30% and maintaining low latency up to 256K tokens. In retrieval-heavy settings, it improves retrieval accuracy by over 20% compared to prior methods, effectively identifying critical information in long inputs.
@@
Future work could explore several promising directions. First, relying on punctuation for sentence segmentation can lead to inaccurate boundaries, especially in texts with sparse or irregular punctuation, which harms semantic coherence. Using NLP-based segmentation tools, such as those in the Natural Language Toolkit Bird (2006), could improve accuracy. Second, the fixed semantic factor may reduce adaptability to texts with varying sentence lengths and structures. Third, many recent reasoning models generate chain-of-thought (CoT) traces with each sentence representing a self-contained logical step, making SentenceKV a natural fit. Future evaluations could target dedicated CoT datasets such as DeepSeek-R1 Guo et al. (2025) and GSM-Infinite Zhou et al. (2025) to assess this potential. Lastly, incorporating lightweight on-device similarity search or mixed-precision sentence embeddings could reduce CPU–GPU round-trip overhead under long-context scenarios.
Acknowledgments
We thank the anonymous reviewers for their valuable suggestions, particularly on evaluating additional benchmarks, as well as their constructive comments on analyzing the efficiency of our method. We also thank Zhenyu Liu from the Department of Electrical, Computer, and Systems Engineering at Rensselaer Polytechnic Institute for his guidance on the implementation details.
Appendix A Appendix
@@ A.2 SentenceKV: pseudocode
0: Prompt tokens, token budget $\tau$, semantic keeping factor $r$, observation window size NN
@@
9:
10:Decoding Phase:
11: Initialize an empty sentence cache $Q_{s}$
12:for $t=1,2,\dots$ do
13:  Generate next token $x_{t}$ and compute query $q_{t}$
14:  Append $q_{t}$ to $Q_{s}$
@@
20:  if a sentence boundary is reached then
21:   Reset $Q_{s}$ for the next sentence
22:  end if
23:end for
@@ A.3 Latency and Memory Footprint
For the Longchat-v1.5-7B LMSYS (2024) model, SentenceKV similarly demonstrates considerable efficiency gains. The peak memory footprint at the maximum context length (256k) is reduced from the OOM (out-of-memory) scenario encountered by Full KV caching to a manageable 48.31 GB. Latency also remains stable around 16.9 ms, whereas Full KV caching becomes infeasible beyond 64k tokens due to excessive memory demands. The details of the results are shown in the Appendix Table 4 and Table 5.
@@
A.4 Quest Sensitivity to Chunk Size
The Quest Tang et al. (2024) cache‐offloading baseline divides the context into fixed‐length chunks and retains only a subset per chunk. Prior work typically uses a chunk size of $8$ tokens, but this granularity differs markedly from SentenceKV, where one bucket naturally corresponds to a full sentence ($\sim\!25\text{--}30$ tokens on our corpora). To ensure a fair comparison on NIAH Kamradt (2023), we therefore re‐evaluated Quest with larger chunk sizes that better match the average sentence length.
Method Quest (Chunk Size 32) Quest (Chunk Size 16) SentenceKV (Adaptive)
NIAH Acc. % 48.3 96.1 97.5
Table 6: Quest results on NIAH under different settings. Model: Llama-3.1-8B with a token budget $\tau=128$.
Quest proves highly sensitive to chunk granularity: reducing the chunk size from $32$ to $16$ tokens nearly doubles accuracy (Table 6). In contrast, SentenceKV employs adaptive buckets (median length $25\text{--}30$ tokens) and attains 97.5 % accuracy with no manual tuning, highlighting the robustness of sentence-aligned grouping.
A.5 Efficiency/Accuracy Trade Off
[…]
SnapKV SentenceKV Increase
Latency (ms/token) 15 16 6%
NIAH Acc. (%) 78.2 97.5 +25%
Table 7: Trade-off on NIAH ($\tau=128,\ r=3$).
Fixing $\tau{=}128$ and semantic-keeping factor r=3r{=}3 on NIAH, SnapKV decodes in $15\;\text{ms}$ per token; SentenceKV needs only $1\;\text{ms}$ extra yet boosts accuracy by 25% (Table 7). Such minor latency growth is often preferable to the sharp accuracy drop that accompanies more aggressive compression.
A.6 GPU Profiling Details
To illustrate how SentenceKV achieves reductions in both memory and latency, we present a fine-grained GPU profile showing (i) the set of KV cache blocks retained in GPU memory during decoding, and (ii) the breakdown of decoding time at a 64K-token context length.
Memory composition at 256 K context (Llama-3.1-8B).
Peak usage reported by torch.cuda.max_memory_allocated() drops from 89.71 GB (Full KV) to 52.71 GB with SentenceKV. The residual footprint consists of:
  1. (1)
    KV cache (subset only): 33.55 GB $\rightarrow$ 0.10 GB after retrieving just $\sim\!1\text{K}$ top tokens per head.
  2. (2)
    Sentence-level semantic vectors: tens of MB; negligible relative to the model.
  3. (3)
    Temporary tensors: short-lived attention buffers, now smaller because attention is computed on far fewer keys/values.
  4. (4)
    Model weights: feed-forward and projection matrices dominate the remaining 50+ GB; they stay on-GPU for all methods.
The reported peak includes allocator fragmentation and transient buffers, yet the relative 37 GB reduction clearly evidences SentenceKV’s compression efficacy.
Latency breakdown at 64 K context (RULER niah_single).
Profiling on an H100-96 GB NVL with $\tau=1024,\ \kappa=2$ yields:
  1. (1)
    Prefill bucket construction (one-time): 4 s for sentence grouping and top-$k$ token selection.
  2. (2)
[…]
  3. (3)
    Sentence ranking (per step): $0.050$ s (CPU dense similarity search).
Even if the $0.073$ s per-step overhead is conservatively added to the computation latency shown in Figure 5, SentenceKV remains markedly faster than Full KV because drastically fewer keys and values participate in attention.
[…]
A.7 Effect of Sentence Length
SentenceKV ranks sentences by relevance until the remaining token budget $\tau$ is exhausted. One possible drawback is that an unusually long but highly relevant sentence could consume a disproportionate share of the budget, thereby excluding other useful context.
To address this, we apply SnapKV-style Li et al. (2025) prefiltering globally across all input tokens during the prefill stage. Specifically, instead of selecting top tokens from each sentence individually, SnapKV selects the top $r\cdot\tau$ most important tokens from the entire document, regardless of sentence boundaries. As a result, the number of retained tokens per sentence is determined by its relative semantic contribution, rather than its raw length. Even if a sentence is long, it will retain fewer tokens if it is not highly informative.
In practice, this global filtering strategy significantly reduces token retention from long but unimportant sentences. Our empirical analysis of LongBench shows that sentence lengths are generally concentrated around the mean, with very long sentences being rare. After filtering, token counts per sentence are typically modest and broadly distributed, helping to avoid scenarios where a single sentence dominates the token budget.
Future work could include detecting abnormally long retained token spans (“buckets”) during prefill and applying simple mitigation strategies. For example, one could compute the mean and standard deviation of sentence lengths in the input, and if a sentence exceeds $\text{mean}+n\times\text{std}$, split it into smaller sub-spans. Such a lightweight check would add minimal computational overhead while preventing outlier buckets from skewing the retrieval process.